OpenStudy (jackiem272):

the human intestine contains Escherichia coli bacteria. in the laboratory this bacteria grows exponentially with a relatively growth of k=0.41 per hour. assume that 1,000 bacterial cells are present at time t=0 a) Find a formula for the number P(t) of bacteria at time t. b)How many bacteria cells will there be 5 hours from now? c) When will there be 10,000 cells

OpenStudy (alekos):

Standard equation is usually P(t) = P(0)e^kt where P(0) = 1,000 and k = 0.41 (I'm assuming)

OpenStudy (alekos):

To work out b) just substitute P(5) in the equation

OpenStudy (alekos):

For c) substitute P(t) = 10,000 and then make t the subject of the formula